### v3

# Quantum Multimodal Contrastive Learning Framework

Abstract
In this paper, we propose a novel framework for multimodal contrastive learning utilizing a quantum encoder to integrate EEG (electroencephalogram) and image data. This groundbreaking attempt explores the integration of quantum encoders within the traditional multimodal learning framework. By leveraging the unique properties of quantum computing, our method enhances the representation learning capabilities, providing a robust framework for analyzing time series and visual information concurrently. We demonstrate that the quantum encoder effectively captures intricate patterns within EEG signals and image features, facilitating improved contrastive learning across modalities. This work opens new avenues for integrating quantum computing with multimodal data analysis, particularly in applications requiring simultaneous interpretation of temporal and visual data.
1 Introduction
Knowing how the human brain processes visual information is a challenging task that connects neuroscience and artificial intelligence (AI). The brain uses a complex system of neural activities across different regions to analyze what we see. This understanding has helped in creating deep learning methods, like convolutional neural networks (CNNs). Further research into how the brain processes visual information in real-life situations could drive new AI advancements.
Recent studies have improved our knowledge by observing brain activity through various techniques. Electroencephalography (EEG) is particularly useful because it is non-invasive, portable, and can capture fast changes in brain activity. EEG gives us a unique way to study how the brain processes and recognizes visual information as it happens in real time. Recently, EEG has seen many interesting applications of machine learning, such as predicting treatment outcomes in psychiatry [1], identity recognition [2], and robotic control [3], etc. These applications rely on the clear and real-time temporal information provided by EEG.
In recent years, contrastive learning has emerged as a powerful technique in self-supervised learning, especially for multimodal data. By learning to associate related pairs from different modalities, contrastive learning has shown impressive results in tasks such as image-text alignment and audio-visual correspondence. However, traditional approaches are constrained by classical computing limitations, particularly when dealing with high-dimensional, complex data like EEG. Decoding images from EEG signals presents a promising avenue for studying brain processes, yet it remains difficult due to the low quality and unpredictability of EEG signals. Overcoming these challenges is crucial for advancing our understanding of visual processing and improving EEG-based image decoding and brain-computer interfaces (BCIs).
Earlier attempts to decode images from EEG data were limited by small datasets, which restricted the ability of models to generalize. More recent efforts have utilized larger datasets, often employing methods where images are presented rapidly to elicit brain responses. However, these methods remain constrained by the limitations of classical neural networks. Despite these advancements, current approaches still fall short, underscoring the need for EEG processing techniques that are better suited to the brain’s mechanisms and the unique characteristics of EEG data.
Although quantum methods have not yet outperformed traditional approaches, recent findings in [4] have shown the potential advantages of quantum encoders for EEG processing. Motivated by these findings, we decided to explore the use of quantum techniques in this context. This led us to introduce a Quantum Multimodal Contrastive Learning Framework (QMCL), an early and exploratory attempt to harness the unique properties of quantum computing, such as superposition and entanglement, to improve the learning process across different data modalities. By integrating Variational Quantum Circuits (VQCs) with classical neural networks, our approach can explore a larger solution space, potentially enabling more efficient and effective learning of multimodal representations. While this hybrid quantum-classical framework is still in its early stages, it holds promise for advancing beyond the capabilities of classical methods and opening up new possibilities for applications that were previously out of reach.
Our QMCL framework is an exploratory approach to a relatively challenging task, offering a promising new direction in processing complex data like EEG. This paper contributes the following:
  • To the best of our knowledge, this is the first work to utilize a VQC architecture for self-supervised cross-modal contrastive learning.
  • We propose a Quantum Multimodal Contrastive Learning Framework (QMCL) and apply it to a zero-shot classification task involving EEG and images.
  • We are the first to use a quantum algorithm on one of the largest EEG-image datasets, setting a new benchmark for future quantum contrastive learning research.
  • We specifically propose four hybrid quantum-classical time-series encoders tailored for QMCL and introduce the Quantum-CLIP-ViT image encoder for this task.
2 Related Works
2.1 Classical Contrastive Learning
Contrastive learning has achieved significant success in unimodal settings. Early models such as SimCLR [5] utilize random data augmentations to generate positive sample pairs and employ cosine similarity for unsupervised learning. Another prominent model, MoCo (Momentum Contrast) [6], introduces a momentum encoder and dynamic dictionary updating, effectively handling large-scale data and outperforming SimCLR in performance. BYOL (Bootstrap Your Own Latent) further advances contrastive learning by removing the need for negative samples, leveraging self-supervised learning to generate high-quality feature representations.
In multimodal contrastive learning, CrossCLR [7] stands out as a significant model that extends contrastive learning to cross-modal representation learning. CrossCLR leverages the correlations between different modalities to improve performance under unsupervised learning. Additionally, CLIP (Contrastive Language-Image Pre-training) [8] marks a milestone in multimodal contrastive learning by simultaneously learning correspondences between text and images, achieving strong multimodal understanding in an unsupervised setting. ALIGN (A Large-scale Image and Noisy Text Embedding) is similar to CLIP, aligning images and text within a shared embedding space using large-scale pre-training, demonstrating the practical potential of multimodal contrastive learning.
These models illustrate the evolution of contrastive learning from unimodal to multimodal applications, highlighting its broad applicability across various data modalities. From unimodal models like SimCLR and MoCo to multimodal models like CLIP and ALIGN, contrastive learning has become a key technique in unsupervised learning, driving advancements in numerous fields. Contrastive learning has shown great success in unimodal settings. Early models like SimCLR use random data augmentations to generate positive pairs and cosine similarity for unsupervised learning. MoCo (Momentum Contrast) introduces a momentum encoder and dynamic dictionary, improving performance on large-scale data. BYOL (Bootstrap Your Own Latent) advances the field by removing the need for negative samples, relying on self-supervised learning for high-quality feature representations.
2.2 Multimodal Contrastive Learning
In recent years, after the success of the traditional contrastive learning models on the same modal data like text and image [9, 10, 11, 12], the development of multimodal contrastive learning has reached significant advancements in the field of self-supervised learning, particularly in tasks that contain the integration of multiple types of data. This method leverages the strengths of various modalities (e.g., text, images, video) to boost model generalization across diverse datasets. Multimodal contrastive learning aligns representations from different modalities within a shared embedding space, facilitating robust, modality-invariant feature learning. This enhances capabilities in cross-modal retrieval and zero-shot learning. Typically, a two-tower network architecture processes each modality independently, with outputs converging in the embedding space where contrastive loss minimizes distances between similar pairs and maximizes distances between dissimilar ones. One of the most popular and successful multimodal contrastive learning framework is CLIP [13], which project both the image and text in to the same feature space. Nevertheless, because datasets containing both time-series signals like EEG and image data are quite rare, there has been little research applying contrastive learning methods to this combination of temporal and visual information. To our best knowledge, [14] is maybe the first work introduced the EEG-image contrastive learning on obtaining the EEG-image representation for image reconstruction downstream task but do not do the zero-shot classification. [15] introduced the EEGClip network for joint representation learning between EEG signal and image but it just do supervised learning. [16] first try to design the EEG encoder on EEG-image contrastive learning , but the work only modified the encoders. This area remains largely uncharted and calls for new, specialized contrastive learning techniques to handle these joint time-series and image modalities effectively.
Decoding visual information from the human brain has long been a challenging task at the intersection of neuroscience and computer science [17, 18, 19, 20]. Although significant progress has been made in understanding static visual inputs, quickly and accurately extracting meaningful information from natural images remains difficult [21, 22]. Early research primarily relied on functional magnetic resonance imaging (fMRI) [23, 24, 25], which has demonstrated the ability to capture meaningful content and structural details from visual processing in the brain. However, fMRI depends on detecting changes in blood oxygen levels, leading to a time lag of several seconds per stimulus, which limits its effectiveness in real-time applications. Moreover, fMRI is expensive and requires large, stationary equipment.
In contrast, electroencephalography (EEG) offers higher temporal resolution, immediate data feedback, and portable, cost-effective hardware, making it a promising candidate for personal brain-computer interface technology. However, current methods using EEG to extract semantic information for image classification have not achieved satisfactory results [26, 27, 16], highlighting the need for improved approaches. Previous methods often relied on supervised learning techniques with a limited set of image categories, ignoring the inherent correlations between visual stimuli and neural responses [27, 28, 15]. These limitations reduce their effectiveness in real-world scenarios that require the generalization to recognize novel, unfamiliar objects. To address these challenges, [29] first attempted zero-shot classification using the largest available EEG-image database [20] with a multilayer MLP and joint EEG-image-text representation, while [16] employed a contrastive learning method. However, [16] used a basic contrastive learning framework based on CLIP [13]. The recent work by [30] improved upon this by using an enhanced neural encoder and an updated CLIP encoder, achieving better results than [16]. Nevertheless, these approaches remain confined to classical machine learning frameworks, and to our knowledge, no similar attempts have been made in the field of quantum machine learning. Our work builds upon the [30] framework and EEG encoder by introducing a self-supervised learning method for EEG-based image decoding and incorporating quantum layers into both the neural and image encoders. This represents a pioneering effort in this field.
2.3 Quantum Machine Learning on Image Data
The integration of quantum computing into machine learning has opened new field for solving complex computational tasks, particularly in the field of image classification. Quantum machine learning (QML) leverages the principles of quantum mechanics, such as superposition and entanglement, to process information in ways that classical systems cannot achieve efficiently. This section explores the application of QML techniques to image data, focusing on the development and evaluation of hybrid quantum neural networks (QNNs) tailored for image classification tasks. [31] introduces two quantum machine learning models designed to address the computational challenges of image classification posed by the growing volume of visual data. One is a hybrid quantum neural network with parallel quantum circuits, outperforming its classical counterpart while using significantly fewer parameters. The other is the model uses a quanvolutional layer to map the input image into lower-resolution feature maps. Despite having fewer parameters, it matches the performance of a classical convolutional neural network. [32] extending the QNN to a scalable variational quantum deep neural network (VQDNN) that supports end-to-end training. That work designing three classifiers for different image recognition tasks using hybrid PCA-VQDNN, amplitude encoding-based VQDNN, and hybrid encoding-based VQDNN to outperform classical neural networks even with a limited number of qubits available in current hardware, suggesting that the approach is promising for practical applications in image recognition using quantum computing. In the survey [33], several recent studies on image classification using quantum machine learning are mentioned. This research direction, compared to traditional machine learning, is still very young but rapidly growing. The integration of quantum computing into machine learning offers new solutions for complex tasks like image classification. Quantum machine learning (QML) utilizes principles like superposition and entanglement to process information more efficiently than classical systems. This section examines the use of QML techniques in image classification, focusing on hybrid quantum neural networks (QNNs). [31] introduces two QML models: a hybrid QNN with parallel quantum circuits that outperforms its classical equivalent with fewer parameters, and a model using a quanvolutional layer to achieve similar performance to classical CNNs with less complexity. [32] extends QNNs to scalable variational quantum deep neural networks (VQDNNs) with end-to-end training, demonstrating superior performance in image recognition tasks even with limited qubits. The survey [33] highlights the rapid growth of this emerging field, despite being in its early stages compared to traditional machine learning.
2.4 Quantum Machine Learning on Time-Series Data
Similar to image data, there have also been recent attempts to apply quantum machine learning to time-series data. The goal is to leverage the properties of quantum information to either reduce the number of model parameters while maintaining the same performance or achieve better data representations. The hybrid quantum-classical VQC algorithm for time-series, in particular, preserves the traditional neural network’s ability to extract temporal and/or spacial structural features while mapping the learned representations to a Hilbert space via quantum layers for feature learning. [34] presents two quantum-classical hybrid machine learning architectures inspired by the multilayer perceptron (MLP). The models were evaluated on both univariate time series (Mackey-Glass time series and USD-to-euro currency exchange rate) and multivariate time series (Lorenz attractor and Box-Jenkins Gas Furnace). The results were competitive with MLP, CNN, and long short-term memory (LSTM) models with a similar number of trainable parameters, suggesting that integrating quantum computing elements could enhance classical models. [35] introduced quantum computing methods for time-series reconstruction and classification, framing these tasks as a quantum unconstrained binary optimization (QUBO) problem solvable by quantum processors. The approach discretizes time-series data, converting it into a set cover problem, enabling efficient reconstruction and semi-supervised classification. The method shows competitive results with classical techniques, using less data, and offers robustness in handling data permutations. [36] introduce a hybrid quantum-classical LSTM model, termed QLSTM. They demonstrate that QLSTM effectively learns various types of temporal data and, in some cases, converges faster or achieves better accuracy than its classical counterpart. The variational approach used in QLSTM reduces the requirements for qubit counts and circuit depth, making it more suitable for implementation on noisy intermediate-scale quantum (NISQ) devices. For the specific EEG application on QML, [4] provided Quantum-EEGNet (QEEGNet), a novel hybrid neural network that integrates quantum computing with the classical EEGNet [37] architecture to enhance EEG encoding and analysis. By incorporating quantum layers, QEEGNet can capture more intricate patterns in EEG data and offer computational advantages.The findings of that work demonstrate the significant potential of quantum-enhanced neural networks in EEG analysis, pointing to promising new directions for research and practical applications.
Figure 1: Overview of the Multimodal Quantum Contrastive Learning (QMCL) framework. Multiple modal data sources are first processed through classical pre-processing modules. The processed data are then encoded using multimodal quantum encoders, with each encoder corresponding to a specific data modality. The encoded representations are compared using contrastive loss functions to align the features across different modalities.
Figure 2: Application of QMCL to EEG-image recognition. Image embeddings are generated using a frozen CLIP-ViT model, while EEG embeddings are produced by a trainable EEG encoder. Both embeddings are processed through quantum neural network (QNN) layers to compute the similarity of each class, enabling the zero-shot recognition task.
2.5 Quantum Contrastive Learning
Recent advances in quantum machine learning have focused primarily on unimodal data, exploring quantum-enhanced methods for tasks like image classification and time series prediction. All the works mentioned below concentrate on single-modality data, without addressing the complexities of learning across multiple modalities. [38] introduces a hybrid quantum-classical architecture for self-supervised learning of visual representations. The study reveals a numerical advantage for small-scale QNNs in learning visual representations, and the quantum models perform comparably to classical models on real quantum devices. This paper [39] proposes a model that fuses Supervised Contrastive Learning (SCL) with VQC and employs Principal Component Analysis (PCA) for dimensionality reduction. The approach effectively addresses the challenge of limited training data and demonstrates potential applications in medical image analysis. Experimental results show that the model achieves high accuracy across several medical imaging datasets, especially when using a small number of qubits (e.g., 2 qubits), highlighting the advantages of quantum computing. [40] presents a novel quantum-enhanced supervised contrastive learning model focused on addressing data scarcity issues. Experiments demonstrate that this model achieves significant accuracy in image classification tasks, even with very limited labeled data. The model exhibits stable performance on real quantum devices, proving its adaptability when faced with limited data. This paper [41] proposes a quantum-enhanced self-supervised learning framework for effective mental health monitoring. The framework uses a quantum-enhanced LSTM encoder to improve representation learning for time series data through contrastive learning. The study shows that the model significantly outperforms traditional self-supervised learning models in predictive capability and achieves high F1 scores across multiple datasets.
Table 1: Comparison of the top-1 accuracy results of QSTConv, QSTConv-GA, QNervFormer and QNervFormer-GA on the ThingsEEG test Dataset for 200-way zero-shot classification.
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6
QSTConv $3.2\pm 1.5\%$ $2.1\pm 1.9\%$ $4.5\pm 0.7\%$ $3.5\pm 3.4\%$ $2.0\pm 2.5\%$ $1.8\pm 1.2\%$
QSTConv-GA $4.2\pm 1.4\%$ $3.0\pm 1.0\%$ $3.3\pm 1.3\%$ $6.0\pm 1.2\%$ $2.6\pm 1.4\%$ $3.6\pm 1.1\%$
QNervFormer 4.3±$\pm 1.7\%$ $4.1\pm 1.4\%$ $\textbf{5.6}\pm 1.5\%$ $5.6\pm 1.4\%$ $4.3\pm 0.4\%$ $4.4\pm 0.7\%$
[…]
Subject 7 Subject 8 Subject 9 Subject 10 Average
QSTConv $3.4\pm 2.1\%$ $4.6\pm 1.1\%$ $3.8\pm 1.5\%$ $6.0\pm 0.7\%$ $3.5\pm 1.7\%$
QSTConv-GA $4.9\pm 0.5\%$ $4.4\pm 0.7\%$ $3.9\pm 1.1\%$ $3.9\pm 1.2\%$ $4.0\pm 2.5\%$
QNervFormer $5.5\pm 1.4\%$ $4.8\pm 2.1\%$ $\textbf{5.9}\pm 2.0\%$ $\textbf{6.6}\pm 1.4\%$ $5.1\pm 1.4\%$
QNervFormer-GA $\textbf{6.4}\pm 1.2\%$ $\textbf{5.8}\pm 1.3\%$ $5.6\pm 1.7\%$ $6.1\pm 1.9\%$ $\textbf{5.2}\pm 1.3\%$
Table 2: Comparison of the top-5 accuracy results of QSTConv, QSTConv-GA, QNervFormer and QNervFormer-GA on the ThingsEEG test Dataset for 200-way zero-shot classification.
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6
QSTConv $12.6\pm 4.8\%$ $9.6\pm 6.1\%$ $16.8\pm 1.4\%$ $12.4\pm 11.0\%$ $7.6\pm 6.1\%$ $8.9\pm 5.6\%$
QSTConv-GA $16.7\pm 5.3\%$ $13.4\pm 1.9\%$ $15.0\pm 2.1\%$ $\textbf{24.4}\pm 2.1\%$ $11.4\pm 2.2\%$ $16.7\pm 2.7\%$
QNervFormer $\textbf{18.0}\pm 3.7\%$ $\textbf{18.6}\pm 2.2\%$ $\textbf{21.9}\pm 0.8\%$ $23.1\pm 1.4\%$ $\textbf{15.5}\pm 1.1\%$ $\textbf{20.2}\pm 2.9\%$
QNervFormer-GA $17.4\pm 2.9\%$ $17.7\pm 2.3\%$ $20.2\pm 2.1\%$ $22.4\pm 1.9\%$ $14.5\pm 1.8\%$ $19.5\pm 2.2\%$
Subject 7 Subject 8 Subject 9 Subject 10 Average
QSTConv $13.3\pm 6.0\%$ $20.0\pm 2.4\%$ $16.3\pm 3.3\%$ $19.7\pm 3.9\%$ $13.7\pm 5.0\%$
QSTConv-GA $16.6\pm 2.3\%$ $20.0\pm 2.1\%$ $13.5\pm 1.6\%$ $20.3\pm 2.7\%$ $16.8\pm 2.5\%$
QNervFormer $20.6\pm 2.4\%$ $19.8\pm 8.5\%$ $\textbf{20.3}\pm 5.4\%$ $23.7\pm 1.4\%$ $20.2\pm 3.0\%$
QNervFormer-GA $\textbf{23.0}\pm 2.9\%$ $\textbf{24.5}\pm 2.3\%$ $19.3\pm 3.2\%$ $\textbf{24.8}\pm 2.1\%$ $\textbf{20.3}\pm 2.4\%$
The proposed Quantum Multimodal Contrastive Learning (QMCL) framework is designed to leverage the advantages of quantum computing for learning rich and complex representations across multiple data modalities. Unlike previous quantum-enhanced contrastive learning methods that focus exclusively on a single modality, QMCL introduces a hybrid quantum-classical architecture capable of handling and integrating data from diverse sources such as images, text, and time series. As the application example of QMCL in this work, we propose a QMCL framework that integrates various quantum-enhanced EEG encoders with a quantum CLIP-ViT for multimodal contrastive learning. The core idea is to leverage the power of quantum computing to encode both EEG signals and images into high-dimensional quantum states and perform contrastive learning between these modalities. The architecture is designed to maximize the similarities between corresponding EEG-image pairs while pushing apart non-corresponding pairs, thus enabling effective multimodal representation learning. The whole architecture illustrated in Fig 2.
2.6 Image Processing with QCLIP-ViT
For image data, we employ the Quantum CLIP-ViT (QCLIP-ViT). The image is first divided into patches, which are flattened and linearly projected. Positional embeddings are added to preserve spatial information, followed by a series of transformer layers, including multi-head self-attention and feed-forward layers. The processed image embeddings are then passed through quantum encoding layers, which map the classical embeddings into quantum states, preparing them for contrastive learning with the EEG embeddings.
Figure 3: The architecture of Quantum-CLIP-Vision transformer (QCLIP-ViT), it uses the pretrained CLIP-ViT weight and project it into qubits’ embedding sopace.
Algorithm 1 Quantum Multimodal Contrastive Learning Framework (QMCL)
1:Input: (Image, EEG) $\triangleright$ stimulus & response
2:Model: $Enc_{img}$: QCLIP-ViT or its variance, $Enc_{eeg}$: QSTConv, QNervFormer or its variance
3:
4:# E : $(batch,channel,electrode,data\ sample)$triangleright batch of input EEGs
5:# I : $(batch,channel,height,width)$triangleright batch of input images
6:
7:# $\tau$ : learned temperature parameter
8:
9:# extract normalized representations from the raw image and EEG
10:$E_{f}$ = Norm(Linear($Enc_{eeg}$(E)))
11:$I_{f}$ = Norm($Enc_{img}$(I)) $\triangleright$ can be obtained before training
12:
13:
14:# scaled pairwise cosine similarity
15:logits = dot($E_{f}$, $I_{f}$.t) $\times$ $e^{\tau}$
16:
17:# symmetric loss function
18:labels = arange(batch) $\triangleright$ self-supervised learning label
19:$loss_{e}$ = CrossEntropyLoss(logits, labels, axis=0)
20:$loss_{i}$ = CrossEntropyLoss(logits, labels, axis=1)
21:$total\_loss$ = ($loss_{e}$ + $loss_{i}$) / 2
2.7 EEG Processing with Quantum Encoders
We implement several variations of quantum encoders tailored for EEG data which are inspired from our previous work [30], each designed to capture different aspects of the time-series signals:
2.7.1 QNervFormer
Combines spatial and temporal convolutional layers followed by multi-head self-attention and cross-attention layers, ultimately passing the features through quantum encoding layers. The model shows in Fig 4.
2.7.2 QNervFormer-GA
Extends QNervFormer by incorporating graph attention mechanisms [42], enhancing the spatial relationships between EEG channels. The model shows in Fig 5.
2.7.3 QSTConv
Inspired by EEGNet, the model first do spatial than do temporal convolutions, followed by quantum encoding layers.
2.7.4 QSTConv-GA
Utilize spatial and temporal convolutions, with the latter also incorporating graph attention layers, followed by quantum encoding layers. The QSTConv series models are illustrated in Fig 6. Each EEG encoder processes the time-series data through classical pre-processing modules, which include convolutional and attention layers, before encoding the output into quantum states. These quantum-encoded EEG features are then aligned with the image embeddings for contrastive learning.
Figure 4: Architecture of the QNervFormer model for time-series data encoding. The model processes time-series data through spatial and temporal convolutions, followed by multi-headed self-attention mechanisms. The final quantum encoding layers enhance the representation of the input data for downstream tasks. Where IN and LN note instance and layer normalization, respectively.
Figure 5: Architecture of QNervFormer-GA, an extension of QNervFormer with integrated graph attention.
Figure 6: Comparison of QSTConv and QSTConv-GA architectures for time-series data encoding. The QSTConv model applies spatial and temporal convolutions with batch normalization (BN) and ELU activation, followed by quantum encoding layers. The QSTConv-GA model introduces graph attention as an additional feature extraction mechanism, integrating it into the pre-processing pipeline before quantum encoding.
This superposition allows the quantum encoding layer to encode and process more information than a classical layer of comparable size.
2.7.5 Quantum Encoding Layer
The quantum encoding layer is to integrate quantum computing capabilities into classical neural networks. By incorporating a quantum circuit within a classical deep learning model, the quantum encoding layer aims to leverage the unique properties of quantum mechanics, such as superposition and entanglement, to enhance the performance and capabilities of machine learning models. It leverages a parameterized quantum circuit defined by mapping classical EEG features as a set of qubits $\mathbf{x}\in\mathbb{R}^{n_{\text{qubits}}}$ and has trainable weights $\mathbf{w}\in\mathbb{R}^{n_{\text{layers}}\times n_{\text{qubits}}}$. The first step in the quantum encoding layer is encoding the input. Each qubit $q_{i}$ is rotated based on the input data $x_{i}$ using the rotation gate RY (rotation-Y gate):
$RY(x_{i})=\exp\left(-i\frac{x_{i}}{2}\sigma_{y}\right),$ (1)
where $\sigma_{y}$ is the Pauli-Y matrix:
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (2)
Second, there are some parameterized layers in the quantum encoding layer module, in this QEEGNet model, we use a ring pattern of CNOT (controlled-NOT) gates. The CNOT gate is a two-qubit gate where the state of one qubit (the control qubit) determines whether to flip the state of another qubit (the target qubit). The CNOT gate is a two-qubit quantum gate represented by the following 4x4 unitary matrix:
${CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (3)
A ring pattern involves applying CNOT gates such that each qubit is entangled with its neighbor, and the last qubit is entangled with the first qubit, forming a closed loop or "ring." This pattern ensures that entanglement is spread throughout the entire set of qubits. For each layer $l$ in $n_{\text{layers}}$, a ring pattern of CNOT gates entangles the qubits:
$CNOT(q_{i},q_{(i+1)\mod n_{\text{qubits}}})\quad\text{for}\quad i=0,1,\ldots,n_{\text{qubits}}-1.$ (4)
Each qubit $q_{i}$ undergoes an additional rotation based on the trainable weight $w_{l,i}$:
$RY(w_{l,i})=\exp\left(-i\frac{w_{l,i}}{2}\sigma_{y}\right).$ (5)
The final stage of the quantum encoding layer module is measurement the states of the qubits. The expectation value of the Pauli-Z operator $\sigma_{z}$ is measured for each qubit, yielding the output:
$\langle\sigma_{z}^{i}\rangle=\langle 0|U^{\dagger}\sigma_{z}^{i}U|0\rangle,$ (6)
where $U$ is the unitary operation representing the quantum circuit. The Pauli-Z operator, also known as the Pauli-Z matrix, is represented by the following 2x2 matrix:
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (7)
This operator is used to measure the z-component of the spin of a qubit and has eigenvalues of +1 and -1, corresponding to the basis states $|0\rangle$ and $|1\rangle$, respectively. The quantum encoding layer integrates quantum computing capabilities into classical neural networks. By embedding a quantum circuit within a classical deep learning model, the quantum encoding layer leverages the unique properties of quantum mechanics—such as superposition and entanglement—to enhance the performance and capabilities of machine learning models. This layer utilizes a parameterized quantum circuit, where classical EEG features are mapped to a set of qubits $\mathbf{x}\in\mathbb{R}^{n_{\text{qubits}}}$, with trainable weights $\mathbf{w}\in\mathbb{R}^{n_{\text{layers}}\times n_{\text{qubits}}}$.
1. Input Encoding: Each qubit $q_{i}$ is rotated based on the input data $x_{i}$ using the rotation-Y (RY) gate:
$RY(x_{i})=\exp\left(-i\frac{x_{i}}{2}\sigma_{y}\right),$ (8)
where $\sigma_{y}$ is the Pauli-Y matrix:
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (9)
2. Parameterized Layers: The quantum encoding layer uses a ring pattern of controlled-NOT (CNOT) gates. A CNOT gate, represented by the following $4\times 4$ unitary matrix, flips the state of a target qubit based on the state of a control qubit:
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (10)
The ring pattern involves applying CNOT gates to entangle each qubit with its neighbor, forming a closed loop. For each layer $l$ in $n_{\text{layers}}$, the CNOT gates are applied as follows:
$\text{CNOT}(q_{i},q_{(i+1)\mod n_{\text{qubits}}})\quad\text{for}\quad i=0,1,\ldots,n_{\text{qubits}}-1.$ (11)
After entanglement, each qubit $q_{i}$ undergoes an additional rotation based on the trainable weight $w_{l,i}$:
$RY(w_{l,i})=\exp\left(-i\frac{w_{l,i}}{2}\sigma_{y}\right).$ (12)
3. Measurement: The final stage involves measuring the states of the qubits. The expectation value of the Pauli-Z operator $\sigma_{z}$ is measured for each qubit, producing the output:
$\langle\sigma_{z}^{i}\rangle=\langle 0|U^{\dagger}\sigma_{z}^{i}U|0\rangle,$ (13)
where $U$ represents the unitary operation of the quantum circuit. The Pauli-Z operator, represented by the matrix:
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ (14)
is used to measure the z-component of the qubit’s spin, with eigenvalues $+1$ and $-1$ corresponding to the basis states $|0\rangle$ and $|1\rangle$, respectively.
2.7.6 Quantum Multimodal Contrastive Learning
The multimodal data from the EEG and image encoders are fed into a contrastive learning framework. We use a cross-modal quantum circuit to entangle the quantum states of the two modalities. The similarity between corresponding EEG-image pairs is maximized while non-matching pairs are pushed apart. The framework uses a contrastive loss function that enforces these relationships across the quantum-encoded embeddings, ensuring that the learned representations capture the inherent multimodal correlations. The training process is conducted using a hybrid quantum-classical approach. The parameters of both the classical pre-processing modules and the quantum circuits are optimized jointly. We employ the parameter-shift rule for the quantum layers, while the classical parts are trained using gradient descent. The training objective is to minimize the contrastive loss, ensuring that the model learns to map corresponding EEG and image data to nearby points in the shared quantum embedding space.
3 Result and Conclusion
3.1 Datasets, Preprocessing and Experiment Details
The ThingsEEG dataset [20] contains extensive EEG recordings collected using a rapid serial visual presentation (RSVP) paradigm, capturing responses from 10 participants to 16,740 natural images from the THINGS database [43]. The dataset includes 1,654 training classes, each with 10 images, and 200 test classes, each with 1 image for zero-shot classification. EEG recordings were performed using 64-channel EASYCAP equipment, with data preprocessed by segmenting trials from 0 to 1,000 ms post-stimulus onset and applying baseline correction using the pre-stimulus mean. EEG responses for each image were averaged across repetitions. The images were resized to 224×224 pixels and normalized before further processing. All the subjects are trained by 200 epochs, we employ the weighted Adam optimizer, setting the learning rate to 0.0002 with $\beta_{1}$ set to 0.5 and $\beta_{2}$ to 0.999. The contrastive learning parameter $\tau$ is initialized to $\log(1/0.07)$. We use 10 qubits and 4 QNN layers for encoding. The results are averaged over 5 training runs using different random seeds.
3.2 Result and Conclusion
The experiment results are shown in TABLE 1 and TABLE 2. The experimental results demonstrate that the QNervFormer-GA model outperforms the other models in both top-1 and top-5 validation accuracy across the 10 subjects in the ThingsEEG dataset. Specifically, QNervFormer-GA achieved the highest average top-1 accuracy at 5.2%, closely followed by QNervFormer at 5.1%. QSTConv-GA and QSTConv lagged behind with average top-1 accuracies of 4.0% and 3.5%, respectively. In terms of top-5 accuracy, QNervFormer-GA again led with an average of 20.3%, slightly ahead of QNervFormer at 20.2%. QSTConv-GA recorded an average of 16.8%, while QSTConv was the lowest at 13.7%. These results highlight the superior generalization ability of QNervFormer-GA, making it the most effective model among those tested for EEG-based image recognition on the ThingsEEG dataset.